# Toward Green and Human-Like Artificial Intelligence: A Complete Survey on Contemporary Few-Shot Learning Approaches

### v2

@@
Abstract.
Despite deep learning’s widespread success, its data-hungry and computationally expensive nature makes it impractical for many data-constrained real-world applications. Few-Shot Learning (FSL) aims to address these limitations by enabling rapid adaptation to novel learning tasks, seeing significant growth in recent years. This survey provides a comprehensive overview of the field’s latest advancements. Initially, FSL is formally defined, and its relationship with different learning fields is presented. A novel taxonomy is introduced, extending previously proposed ones, and real-world applications in classic and novel fields are described. Finally, recent trends shaping the field, outstanding challenges, and promising future research directions are discussed.
Few-Shot Learning, Meta-Learning, Transfer Learning, Learning with Few Samples, Human-Like Learning, Green AI
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: General and reference Surveys and overviewsccs: Computing methodologies Learning paradigmsccs: Computing methodologies Machine learning approachesccs: Computing methodologies Machine learning algorithms
1. Introduction
In the past decade, there has been significant progress in machine/deep learning fields, mainly due to the abundance of computational resources, the availability of massive datasets, and the development of sophisticated algorithms, like the Transformer architecture (Vaswani et al., 2017). The efficacy of machine and deep learning models has been materialized in diverse fields, including computer vision(He et al., 2016), reinforcement learning(Silver et al., 2017), natural language processing(Brown et al., 2020), and biomedicine(Jumper et al., 2021), while also surpassing human performance in some cases(Silver et al., 2017). A vital aspect of these models, considered the ’golden rule’ of modern deep learning, is their training from scratch on a specific task using a large dataset.
@@ 1. Introduction
Parallel to the development of large-scale models, there is a growing interest in integrating such aspects of human-like learning processes within the existing machine learning paradigm. As a result, to bridge the gap between machine and human learning, Few-Shot Learning (FSL) has emerged as a promising direction. More precisely, FSL aims to develop models that can generalize within a specific task, even with a limited number of training samples available. This is typically achieved by using prior knowledge extracted from similar tasks. Notably, such models can adapt to novel tasks with small datasets, which is difficult for conventional machine learning models without the risk of overfitting. Additionally, they can significantly reduce the need for collecting and preprocessing extensive amounts of data for training while also minimizing the time required for adapting to new data. On a more practical level, FSL can also enable learning of rare cases, which is crucial in applications such as drug discovery(Altae-Tran et al., 2017).
@@
Leveraging prior knowledge to adapt to novel tasks has been explored since the 1980s via self-referential learning(Schmidhuber, 1987) and learning biologically plausible rules that enable adaptation to new tasks(Bengio et al., 1991). Other early approaches have also included object recognition from a single image using Bayesian models(Fei-Fei et al., 2006), learning data augmentations(Miller et al., 2000), and metric learning(Fink, 2004). However, the proliferation of FSL and its influence on various aspects of modern deep learning over the past decade necessitate consolidating knowledge on recent advancements in the field.
This paper aims to address this need by providing a comprehensive and up-to-date survey of the field of FSL. While previous surveys have focused on specific types of methods within the FSL paradigm(Wang et al., 2020b; Song et al., 2023a), we aim to give a broader overview of the different types of learning applicable in this setting. Overall, our contributions can be summarized as follows:
  • We provide a general yet useful formal definition of FSL by extending the empirical risk minimization framework used in standard machine learning.
  • We examine FSL’s connections with relevant learning problems, including established approaches such as meta-learning and transfer learning, as well as newer ones such as self-supervised learning.
@@
  • We propose a novel taxonomy that extends previously proposed ones by including classic approaches, such as meta-learning, transfer learning, and data augmentation, and newer ones, such as their probabilistic extensions, Neural Processes, and hybrid methods combining different learning types. We primarily focus on recent advancements while also including well-established methods.
@@
The remainder of this survey is organized as follows. Section 2 provides an overview of FSL, including its formal mathematical formulation, commonly used datasets, and tangential fields of interest. Section 3 introduces the proposed taxonomy and provides an overview of various recent and standard approaches within each category. Section 4 analyzes FSL’s fields of application. Section 5 analyzes multiple trends that have emerged in recent years and how they have shaped the modern landscape of FSL, while Section 6 discusses less explored directions within the field that are promising for future research. Finally, Section 7 concludes the survey.
2. Background
@@
2.1. Problem Formulation
As previously mentioned, FSL refers to a set of learning problems in which machine learning models aim to adapt and be able to generalize given only a limited number of training examples. Wang et al.(Wang et al., 2020b) provide a more formal definition framing FSL using the same terminology as the one used for defining classic machine learning. In general, given a task T and a performance measure P, a model is said to learn when its performance P on the task T is improved by leveraging experience E related to the task (e.g., training data).
@@ 2.1. Problem Formulation
(1) $\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{(x,y)\sim p(x,y)}[\mathcal{L}(f_{\theta}(x),y)]$
@@
(2) $\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i})\right]$
@@
(3) $\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{T\sim q(T)}[\mathbb{E}_{(x,y)\sim p(x,y)}\mathcal{L}(f_{\theta}(x),y)]$
@@
(4) $\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\left[\frac{1}{M}\sum_{j=1}^{M}\left[\frac{1}{N}\sum_{(x,y)\in D_{j}^{Q}}\mathcal{L}(f_{\theta|D_{j}^{S}}(x),y)\right]\right]$
@@ 2.2. N-way K-shot Formulation
Historically, research around FSL has mainly focused on image classification. In this particular setting, one typically refers to the N-way-K-shot classification(Vinyals et al., 2016), in which each support set contains samples from N different classes, with K different examples in each class. As a result, each support set consists of $|S|=NK$ training samples. Similar formulations can also be implemented for different learning problems. For instance, in few-shot semantic segmentation, each task might consist of various images of the same object to be segmented, and in few-shot regression, each task corresponds to a different output function with only a few input-output pairs provided in the support set.
2.3. Datasets
@@
Various datasets suitable for FSL training and evaluation have been proposed throughout the years, with miniImageNet(Vinyals et al., 2016) and Omniglot(Lake et al., 2015) being the most widely used for few-shot image classification. In particular, miniImageNet consists of 60000 images separated into 100 different classes originating from ImageNet, and Omniglot contains images of 1623 handwritten characters, with 20 different samples included for each character. Other commonly used datasets include tieredImageNet(Ren et al., 2018), which alleviates miniImageNet’s lack of class diversity, having 608 classes grouped in 34 class categories, and CIFAR-FS(Bertinetto et al., 2019). Finally, datasets such as CUB-200-2011, tinyImageNet, Aircraft and Stanford Cars, although not initially developed for FSL purposes, have also been used due to their small size and variety of classes.
@@
However, these datasets lack variability in their number of classes and samples per class, and they are also not suitable for evaluating out-of-domain generalization. To combat these issues, datasets such as Meta-Dataset(Triantafillou et al., 2020) and ORBIT(Massiceti et al., 2021) have emerged that combine images from different datasets. Extending this framework, there has also been a bloom of other datasets in recent years, aiming to include images from various domains, such as BSCD-FSL(Guo et al., 2020), which includes agriculture, satellite, and dermatology. Other recent approaches include Meta-Album(Ullah et al., 2022) focusing on a more realistic cross-domain evaluation of FSL models, spanning across ten distinct domains and HARD-MD++(Basu et al., 2022) which is the amalgamation of various tasks considered hard to solve from Meta-Dataset and other FSL benchmarks, spanning across four domains. Finally, Meta Omnium(Bohdal et al., 2023) extends the FSL framework outside of image classification by providing a benchmark for evaluation of generalization across different learning problems such as key-point localization, semantic segmentation, and regression.
@@
Apart from image classification, various other datasets related to specific problems such as semantic segmentation(Li et al., 2020), regression(Gao et al., 2022), pose estimation(Xu et al., 2022a), drug discovery(Stanley et al., 2021) and robotics (Yu et al., 2020) have been proposed, however they usually lack the diversity displayed by modern image classification few-shot benchmarks.
2.4. Relevant Learning Problems
@@
Learning efficiently with few data, transferring knowledge across different datasets, and leveraging auxiliary tasks to enhance learning have been well-known research topics within the machine learning community for years. Under these settings, which are particularly challenging for classic supervised learning, various approaches have been proposed throughout the years. In this section, we compare these types of learning with the FSL paradigm, delineating their similarities and differences.
Meta-Learning
@@
The main goal of meta-learning is to enable faster adaptation to novel tasks by utilizing previously acquired knowledge. To achieve this, meta-learning models extract the common structure from a pool of tasks and use it as an inductive bias for rapid adaptation to unseen tasks with scarce data. In general, meta-learning has been widely used to solve FSL tasks (see Section 3.1). However, its goal of generalizing to novel tasks is not necessarily restricted to the assumption of limited data as in FSL.
Weakly Supervised Learning
@@
This type of learning includes incomplete supervision, where both labeled and unlabeled data are utilized during training, inexact supervision that focuses on training with data that have coarser labels than the ones the model wants to predict, and inaccurate supervision that deals with noisy and incorrect data labels. Although both FSL and weakly supervised learning (WSL) rely on introducing inductive biases by leveraging other data sources, in WSL, these originate from unlabelled data, while in FSL, they usually originate from labeled data from different tasks or pretrained models.
Semi-Supervised Learning
@@
A typical case of weakly-supervised learning is semi-supervised learning, where the model is trained on a few labeled data and a large number of unlabelled data. Typically, the semi-supervised model aims to infer a general structure, acting as an inductive bias, from the unlabelled data and then leverage it to boost its performance on the few labeled data. This bears great similarity with FSL approaches that infer inductive biases from outer sources. However, FSL uses labeled data from other tasks to extract prior knowledge.
Self-Supervised Learning
@@
Similarly to FSL, self-supervised models are initially trained on pretext tasks. However, in self-supervised learning (SSL), these tasks are typically self-generated and consist of unlabelled data, while FSL tasks are labeled and predefined. Additionally, in SSL, there is no restriction on whether these downstream tasks contain only a few data. Generally, the two paradigms have a close relationship since they aim to extract robust and transferable representations that can be used in downstream tasks. As a result, various recent approaches have incorporated self-supervised techniques such as contrastive learning within the FSL framework (see Section 3.5).
Transfer Learning
@@
The goal of transfer learning is to transfer knowledge from a source domain where data is widely available to a target domain where data is limited. Regarding FSL, various transfer learning approaches based on model finetuning have successfully been used (see Section 3.2). However, transfer learning deals with the source (abundant data) and target (limited data) tasks only, and its goal is to transfer knowledge from the former to the latter. On the other hand, FSL deals with multiple tasks, each containing only a few data, and its goal is to extract knowledge from the underlying structure shared among tasks.
Domain Adaptation/Generalization
@@
Domain adaptation aims to train models capable of maintaining a solid performance when evaluated on out-of-distribution samples, while domain generalization extends this framework to create models that exhibit robustness under any distribution shift. Recently, domain adaptation/generalization has been studied within the FSL framework, expanding it from creating models capable of adapting to new tasks with only a few data to also being robust to out-of-distribution tasks and data.
Multi-Task Learning
@@
In multi-task learning (MTL), the goal is to train a model that is jointly optimized on a set of tasks. Each task contributes to the optimization of all tasks, and consequently, performance is increased compared to training a separate model for each task. This type of training is similar to FSL, where shared information is extracted from training tasks and is subsequently used to improve performance on novel tasks. However, MTL focuses on a predetermined set of tasks, while FSL aims to generalize to novel tasks not seen during training.
Zero-Shot Learning
@@
The major difference between zero-shot learning (ZSL) and FSL is the number of available training samples within each task. In particular, FSL tasks contain a small number of training samples, while ZSL contains none. As a result, while FSL relies on adapting to these few samples, ZSL leverages external data, usually from different modalities, to enable learning in these tasks. Overall, both approaches introduce inductive biases that facilitate effective learning in novel tasks, yet their sources differ significantly.
Continual Learning
@@
In continual learning, the goal is to develop models that can incrementally acquire new knowledge from different tasks without their performance deteriorating due to catastrophic forgetting. While both FSL and continual learning incorporate learning across various tasks, FSL methods are trained offline, while continual learning is heavily based on online learning. This shifts the focus of continual learning towards incremental accumulation of knowledge, while FSL focuses on leveraging prior information to generalize within tasks with a limited number of data efficiently.
3. Few-Shot Learning Models

### v3

@@
Abstract
Despite the widespread success of deep learning, its intense requirements for vast amounts of data and extensive training make it impractical for various real-world applications where data is scarce. In recent years, Few-Shot Learning (FSL) has emerged as a learning paradigm that aims to address these limitations by leveraging prior knowledge to enable rapid adaptation to novel learning tasks. Due to its properties that highly complement deep learning’s data-intensive needs, FSL has seen significant growth in the past few years. This survey provides a comprehensive overview of both well-established methods as well as recent advancements in the FSL field. The presented taxonomy extends previously proposed ones by incorporating emerging FSL paradigms, such as in-context learning, along with novel categories within the meta-learning paradigm for FSL, including neural processes and probabilistic meta-learning. Furthermore, a holistic overview of FSL is provided by discussing hybrid FSL approaches that extend FSL beyond the typically examined supervised learning setting. The survey also explores FSL’s diverse applications across various domains. Finally, recent trends shaping the field, outstanding challenges, and promising future research directions are discussed.
keywords:
Few-Shot Learning, Meta-Learning, Transfer Learning, In-Context Learning, Human-Like Learning, Green AI
1 Introduction
In the past decade, there has been significant progress in machine/deep learning fields, mainly due to the abundance of computational resources, the availability of massive datasets, and the development of sophisticated algorithms, like the Transformer architecture [204]. The efficacy of machine and deep learning models has been materialized in diverse fields, including computer vision [75], reinforcement learning [186], natural language processing [19], and biomedicine [96], while also surpassing human performance in some cases [186]. A vital aspect of these models, considered the ’golden rule’ of modern deep learning, is their training from scratch on a specific task using a large dataset.
@@ 1 Introduction
Parallel to the development of large-scale models, there is a growing interest in integrating such aspects of human-like learning processes within the existing machine learning paradigm. As a result, to bridge the gap between machine and human learning, Few-Shot Learning (FSL) has emerged as a promising direction. More precisely, FSL aims to develop models that can generalize within a specific task, even with a limited number of training samples available. This is typically achieved by using prior knowledge extracted from similar tasks. Notably, such models can adapt to novel tasks with small datasets, which is difficult for conventional machine learning models without the risk of overfitting. Additionally, they can significantly reduce the need for collecting and preprocessing extensive amounts of data for training while also minimizing the time required for adapting to new data. On a more practical level, FSL can also enable learning of rare cases, which is crucial in applications such as drug discovery [4].
@@
Leveraging prior knowledge to adapt to novel tasks has been explored since the 1980s via self-referential learning [180] and learning biologically plausible rules that enable adaptation to new tasks [16]. Other early approaches have also included object recognition from a single image using Bayesian models [51], learning data augmentations [143], and metric learning [53]. However, the proliferation of FSL and its influence on various aspects of modern deep learning over the past decade have necessitated consolidating knowledge on recent advancements in the field.
One of the first surveys on FSL was conducted by [219], who provided the first formal definition and taxonomy of existing methods and applications in the field. This taxonomy was later extended in [161] to include various hybrid methods beyond the standard supervised setting. More recently, in [188], a novel taxonomy has been proposed based on how prior information is abstracted in FSL and the different challenges arising in each case. Meanwhile, other surveys such as in [66] have narrowed the focus to meta-learning approaches and their application in the field of FSL. Additionally, several recent surveys have focused on specific application areas of FSL, such as FSL in medical imaging [157], FSL in channel state information-based human sensing [221], and few-shot object detection [110], among others.
However, there has not yet been a comprehensive overview of FSL that covers the full range of modern FSL methods and applications. Specifically, certain existing surveys have primarily focused on meta-learning approaches for FSL [66] while most recent attempts to provide a broader perspective on the field have not captured emerging FSL paradigms such as the trending in-context learning [188, 66]. Meanwhile, discussions regarding the extensions of FSL beyond the supervised learning setting using hybrid methods have been limited in the past [161]. Finally, several surveys [157, 221, 110] have focused exclusively on specific application fields of FSL.
Table 1: Comparison of existing FSL surveys and our survey. Symbol ✓means the topic is covered while ✗ means the topic is not covered.
Survey In-Context Learning Novel Meta- Learning Model Families Non Meta- Learning FSL Methods Hybrid Methods Beyond Supervised FSL
[219]
[161]
[188]
[66]
Ours
This paper aims to address the current lack of a broader overview in the field of FSL by providing a comprehensive and up-to-date survey. In particular, it includes an exploration of novel model families under existing FSL paradigms, such as neural processes and probabilistic meta-learning, incorporates novel FSL paradigms, such as in-context learning, and offers a more detailed review of hybrid FSL methods that extend FSL beyond supervised learning settings. Table 1 illustrates how our survey integrates these various elements extending upon previous surveys in the field. Finally, emerging challenges and future directions for FSL in the era of large-scale deep learning models such as human-like learning and Green AI are discussed. Overall, our contributions can be summarized as follows:
@@
  • We propose a novel taxonomy that builds upon previously proposed ones and extends classic approaches, such as meta-learning, transfer learning, and data augmentation, by introducing new subcategories, including neural processes and probabilistic extensions of meta-learning. We primarily focus on recent advancements while also including well-established methods.
  • We present, for the first time in an FSL-related survey, an analysis of the recently explored in-context learning within the FSL framework, examining its similarities and differences with more traditional FSL approaches as well as various techniques that combine these two paradigms.
  • We explore FSL’s connections with relevant learning problems and review various existing hybrid approaches that extend the standard supervised FSL framework to different settings, including unsupervised FSL, few-shot federated learning, cross-domain FSL, and class-incremental FSL, among others.
@@
The remainder of this survey is organized as follows. Section 2 provides an overview of FSL, including its formal mathematical formulation, commonly used datasets, and related fields of interest. Section 3 introduces the proposed taxonomy and provides an overview of various recent and standard approaches within each category. Section 4 extends the presented taxonomy beyond the supervised FSL setting by introducing hybrid methods at the intersection of FSL and different learning paradigms. Section 5 analyzes FSL’s fields of application. Section 6 analyzes multiple trends that have emerged in recent years and discusses less explored directions within the field that are promising for future research. Finally, Section 7 concludes the survey.
2 Background
@@
2.1 Problem Formulation
As previously mentioned, FSL refers to a set of learning problems in which machine learning models aim to adapt and be able to generalize given only a limited number of training examples. [219] provide a more formal definition framing FSL using the same terminology as the one used for defining standard machine learning. In general, given a task T and a performance measure P, a model is said to learn when its performance P on the task T is improved by leveraging experience E related to the task (e.g., training data).
@@ 2.1 Problem Formulation
$\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{(x,y)\sim p(x,y)}[\mathcal{L}(f_{\theta}(x),y)]$ (1)
@@
$\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i})\right]$ (2)
@@
$\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{T\sim q(T)}[\mathbb{E}_{(x,y)\sim p(x,y)}\mathcal{L}(f_{\theta}(x),y)]$ (3)
@@
$\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\left[\frac{1}{M}\sum_{j=1}^{M}\left[\frac{1}{N}\sum_{(x,y)\in D_{j}^{Q}}\mathcal{L}(f_{\theta|D_{j}^{S}}(x),y)\right]\right]$ (4)
@@ 2.2 N-way K-shot Formulation
Historically, research around FSL has mainly focused on image classification. In this particular setting, one typically refers to the N-way-K-shot classification [205], in which each support set contains samples from N different classes, with K different examples in each class. As a result, each support set consists of $|S|=NK$ training samples. Similar formulations can also be implemented for different learning problems. For instance, in few-shot semantic segmentation, each task might consist of various images of the same object to be segmented, and in few-shot regression, each task corresponds to a different output function with only a few input-output pairs provided in the support set.
2.3 Datasets
@@
Various datasets suitable for FSL training and evaluation have been proposed throughout the years, with miniImageNet [205] and Omniglot [116] being the most widely used for few-shot image classification. In particular, miniImageNet consists of 60000 images separated into 100 different classes originating from ImageNet, and Omniglot contains images of 1623 handwritten characters, with 20 different samples included for each character. Other commonly used datasets include tieredImageNet [174], which alleviates miniImageNet’s lack of class diversity, having 608 classes grouped in 34 class categories, and CIFAR-FS [17]. Finally, datasets such as CUB-200-2011, tinyImageNet, Aircraft and Stanford Cars, although not initially developed for FSL purposes, have also been used due to their small size and variety of classes.
@@
However, these datasets lack variability in their number of classes and samples per class, and they are also not suitable for evaluating out-of-domain generalization. To combat these issues, datasets such as Meta-Dataset [201] and ORBIT [140] have emerged that combine images from different datasets. Extending this framework, there has also been a bloom of other datasets in recent years, aiming to include images from various domains, such as BSCD-FSL [71], which includes agriculture, satellite, and dermatology. Other recent approaches include Meta-Album [203] focusing on a more realistic cross-domain evaluation of FSL models, spanning across ten distinct domains and HARD-MD++ [13] which is the amalgamation of various tasks considered hard to solve from Meta-Dataset and other FSL benchmarks, spanning across four domains. Finally, Meta Omnium [18] extends the FSL framework outside of image classification by providing a benchmark for evaluation of generalization across different learning problems such as key-point localization, semantic segmentation, and regression.
@@
Apart from image classification, various other datasets related to specific problems such as semantic segmentation [125], regression [60], pose estimation [233], drug discovery [190] and robotics [245] have been proposed, however they usually lack the diversity displayed by modern image classification few-shot benchmarks.
2.4 Relevant Learning Problems
@@
Learning efficiently with few data, transferring knowledge across different datasets, and leveraging auxiliary tasks to enhance learning have been well-known research topics within the machine learning community for years. Under these settings, which are particularly challenging for standard supervised learning, various approaches have been proposed throughout the years. In this section, we compare these types of learning with the FSL paradigm, delineating their similarities and differences. Additionally, Section 4 provides a more detailed discussion, where several hybrid methods that lie at the intersection of FSL and these learning paradigms are presented.
[…]
@@
Meta-Learning The main goal of meta-learning is to enable faster adaptation to novel tasks by utilizing previously acquired knowledge. To achieve this, meta-learning models extract the common structure from a pool of tasks and use it as an inductive bias for rapid adaptation to unseen tasks with scarce data. In general, meta-learning has been widely used to solve FSL tasks (see Section 3.1). However, its goal of generalizing to novel tasks is not necessarily restricted to the assumption of limited data as in FSL.
[…]
@@
Weakly Supervised Learning This type of learning includes incomplete supervision, where both labeled and unlabeled data is utilized during training, inexact supervision that focuses on training with data that has coarser labels than the ones the model wants to predict, and inaccurate supervision that deals with noisy and incorrect data labels. Although both FSL and weakly supervised learning (WSL) rely on introducing inductive biases by leveraging other data sources, in WSL, these originate from unlabeled data, while in FSL, they usually originate from labeled data from different tasks or pretrained models.
[…]
@@
Semi-Supervised Learning A typical case of weakly-supervised learning is semi-supervised learning, where the model is trained on a few labeled data and a large number of unlabeled data. Typically, the semi-supervised model aims to infer a general structure, acting as an inductive bias, from the unlabeled data and then leverage it to boost its performance on the few labeled data. This bears great similarity with FSL approaches that infer inductive biases from outer sources. However, FSL uses labeled data from other tasks to extract prior knowledge.
[…]
@@
Self-Supervised Learning Similarly to FSL, self-supervised models are initially trained on pretext tasks. However, in self-supervised learning (SSL), these tasks are typically self-generated and consist of unlabeled data, while FSL tasks are labeled and predefined. Additionally, in SSL, there is no restriction on whether these downstream tasks contain only a few data. Generally, the two paradigms have a close relationship [153] since they aim to extract robust and transferable representations that can be used in downstream tasks. As a result, various recent approaches have incorporated self-supervised techniques such as contrastive learning within the FSL framework (see Section 4.3).
[…]
@@
Transfer Learning The goal of transfer learning is to transfer knowledge from a source domain where data is widely available to a target domain where data is limited. Regarding FSL, various transfer learning approaches based on model finetuning have successfully been used (see Section 3.2). However, transfer learning deals with the source (abundant data) and target (limited data) tasks only, and its goal is to transfer knowledge from the former to the latter. On the other hand, FSL deals with multiple tasks, each containing only a few data, and its goal is to extract knowledge from the underlying structure shared among tasks.
[…]
@@
Domain Adaptation/Generalization Domain adaptation aims to train models capable of maintaining a solid performance when evaluated on out-of-distribution samples, while domain generalization extends this framework to create models that exhibit robustness under any distribution shift [106]. Recently, domain adaptation/generalization has been studied within the FSL framework, expanding it from creating models capable of adapting to new tasks with only a few data to also being robust to out-of-distribution tasks and data.
[…]
@@
Multi-Task Learning In multi-task learning (MTL), the goal is to train a model that is jointly optimized on a set of tasks. Each task contributes to the optimization of all tasks, and consequently, performance is increased compared to training a separate model for each task. This type of training is similar to FSL, where shared information is extracted from training tasks and is subsequently used to improve performance on novel tasks. However, MTL focuses on a predetermined set of tasks, while FSL aims to generalize to novel tasks not seen during training.
[…]
@@
Zero-Shot Learning The major difference between zero-shot learning (ZSL) and FSL is the number of available training samples within each task. In particular, FSL tasks contain a small number of training samples, while ZSL contains none. As a result, while FSL relies on adapting to these few samples, ZSL leverages external data, usually from different modalities, to enable learning in these tasks. Overall, both approaches introduce inductive biases that facilitate effective learning in novel tasks, yet their sources differ significantly.
[…]
@@
Continual Learning In continual learning, the goal is to develop models that can incrementally acquire new knowledge from different tasks without their performance deteriorating due to catastrophic forgetting [105]. While both FSL and continual learning incorporate learning across various tasks, FSL methods are trained offline, while continual learning is heavily based on online learning. This shifts the focus of continual learning towards incremental accumulation of knowledge, while FSL focuses on leveraging prior information to generalize within tasks with a limited number of data efficiently.
3 Few-Shot Learning Models